OpenStudy (anonymous):

Which shows three ordered pairs that are all solutions of the equation y = x + 7? A. (0, 7), (2, 9), and (3, 11) B. (0, 7), (2, 9), and (3, 10) C. (1, 7), (2, 9), and (3, 10) D. (1, 8), (2, 9), and (3, 11)

OpenStudy (.cheesecake.):

Take your coordinates and plug them into the equation to see whether they are true or false

OpenStudy (anonymous):

still don't get it

OpenStudy (alivejeremy):

cause u dumb

OpenStudy (.cheesecake.):

(x,y) plug the numbers from the coordinates in place... 7 = 0 + 7 9 = 2 + 7 11= 3 + 7 are these true or false?

OpenStudy (anonymous):

f

OpenStudy (mrnood):

for instance the first pair x=0, y =7 so in the equation y= x+7 you write 7 =0+7. This is TRUE - so that first point IS a soluiton. You need to do that for all the pairs until you find the option where all 3 are correct

OpenStudy (.cheesecake.):

Good it is indeed false 7= 0+ 7 9= 2+ 7 10 = 3 + 7 true or false?

OpenStudy (anonymous):

then.................

OpenStudy (.cheesecake.):

? are the above statements true or false?

OpenStudy (anonymous):

true

OpenStudy (.cheesecake.):

Good so B would be our answer :)

OpenStudy (anonymous):

thx

OpenStudy (.cheesecake.):

yw